geerky42 (geerky42):

Find the units digit of 13²º¹²

hartnn (hartnn):

3^1= 3^2= 3^3= 3^4=

hartnn (hartnn):

sorry, 13^ ....

geerky42 (geerky42):

13¹ = 3 13² = 169 13³ = 2197 13⁴ = 28561 13⁵ = 371293

hartnn (hartnn):

see the pattern ? 3 9 7 1 3 9 7 1 . .

geerky42 (geerky42):

Well, how can I know what 13²º¹² is?

hartnn (hartnn):

13^(4n) has last digit = 1

geerky42 (geerky42):

So the answer is 1 since 2012 is divisible of 4?

hartnn (hartnn):

yes.

hartnn (hartnn):

any doubts ?